OpenStudy (anonymous):

What is the equation of the line passing through (1, 7) and parallel to the line y = –x + 5 in standard form?

OpenStudy (anonymous):

y=-x+8

OpenStudy (anonymous):

parallel lines have the same slope

OpenStudy (anonymous):

these are the options: x + y = 8 x – y = 8 x + y = –8 x – y = –8

OpenStudy (anonymous):

a

OpenStudy (anonymous):

being parallel to y=-x+5 your required equation is y=-x+k(say) as parallel lines have the same slope since the line passes through (1,7) substitute in the above equation 7=-1+k which gives you k=8 therfore the requied equation is y=-x+8

OpenStudy (anonymous):

so its a

OpenStudy (anonymous):

Solve the following for x and them move -y from the RHS to the LHS.\[\frac{y-7}{x-1}=-1\] \[x+y=8 \]